OpenStudy (anonymous):

Macy wants to buy a skateboard for $42. She has $95 in her account. She spent $12.99 to buy stationary. She also wants to buy some ice cream bars for $3.07 each. What is the maximum number of ice cream bars, n, that Macy can buy so that she has enough money left to buy the skateboard?

OpenStudy (anonymous):

helpppppppp

OpenStudy (danjs):

Total Money in Account = Skateboard + Stationary + 3.07*(Ice Creams n)

OpenStudy (danjs):

Total Money = 95 Skateboard = 42 Stationary = 12.99 Ice Cream = 3.07*n , Where n is the number of ice creams. Put into the equation and solve.

OpenStudy (danjs):

95 = 42 + 12.99 + 3.07*n n = (Round down to closest whole number) of ice creams

OpenStudy (anonymous):

n ≤ 13 n ≥ 13 n ≤ 16 n ≥ 16

OpenStudy (danjs):

95 is < or = to 42 + 12.99 + 3.07n

OpenStudy (anonymous):

why are the question marks still there

OpenStudy (danjs):

reload page, it happens when you paste things

OpenStudy (danjs):

\[95 \le 42 + 12.99 + 3.07*n\]

OpenStudy (anonymous):

that is not one of the answer that is why I am confused

OpenStudy (danjs):

right, that is the equation that the question describes. You need to isolate n, solve for n is less than or equal too......

OpenStudy (danjs):

\[95 \ge 54.99 + 3.07n\] \[95 - 54.99 \ge 3.07n\] \[40.01 \ge 3.07n\] \[\frac{ 40.01 }{ 3.07 } \ge n\] \[n \le 13\]

OpenStudy (danjs):

she can buy up to 13 ice creams and still have enough money for the skateboard